# FrugalPrompt: Reducing Contextual Overhead in Large Language Models via Token Attribution

### v2

@@ 5.4.  Experimental Setup
Attribution
Type
Token
Attribution
Method
$k\%$ Tasks
@@
6.  Result and Analysis
6.1.  Performance across Tasks
Tab. 1 reveals that performance degradation becomes more pronounced as the retention threshold decreases, particularly below 60% token retention, with distinct patterns emerging across tasks. For instance, at 50% retention, sentiment classification accuracy drops by 5-10% and commonsense QA accuracy by 10-15%, while summarization metrics like ROUGE-1 decline by 20-25%, as seen in Figure 2,3. The performance decline is noticeably smaller for higher values of $k$, e.g.,, GlobEnc using $80\%$ tokens achieves close to the baseline performance in classification and QA tasks.
Mathematical reasoning exhibits the steepest fall, with pass@1 scores plummeting to below 15% at lower thresholds, reflecting the loss of critical numerical and logical connectors. We hypothesize that this degradation is more pronounced for reasoning tasks due to their reliance on the sequential integrity of tokens and numerical data. Low $k\%$ thresholds exacerbate this by discarding essential context, amplifying errors in tasks requiring unbroken symbolic chains, while simpler classification and summarization tasks generally tolerate greater omission by inferring from salient cues.
@@ 6.1.  Performance across Tasks
6.2.  Choice of Attribution Method
Following both Tab. 1 and Fig. 2, for classification and QA tasks, GlobEnc outperforms DecompX across all metrics. However, in reasoning, DecompX outshines GlobEnc. DecompX’s subword-level attribution that includes linear activations may help retain more important tokens for reasoning, but might overwhelm the models in other tasks. For summarization (Fig. 3), GlobEnc usually outperforms DecompX, except for the GPT-3.5 model. While DecompX is reported to produce better attribution scores Modarressi et al. (2023), most models find the token sequence filtered by word-level attention, like GlobEnc, to be more useful in retaining performance across most tasks.
6.3.  Performance vs. Cost
The cost per million tokens for each model was taken from the respective provider’s website, at the time of writing, and is reported in Tab. 3. The reader should note that the price is subject to change. The cost of the open-source Llama models was estimated using the lowest available rate on OpenRouter1.
Fig.2, 3 show the performance difference between the reduced set of tokens and 100% tokens for all the models. For sentiment analysis and question answering, we observe that using a reduced set of tokens generally retains the performance and has a softer decline as we go down the cost. However, reducing the tokens sharply declines the performance in reasoning tasks, making it unsuitable for token reduction. The results are mixed for summarization tasks: while most models have a softer performance decline across the metrics, methods such as GPT-3.5 using GlobEnc experience rapid performance degradation.
6.4.  Performance across Models
Figure 2,3 illustrate that for smaller models, token attribution offers negligible cost savings relative to the sharper performance degradation it incurs. Consequently, token reduction proves to be more practical for larger, costlier models such as o3-mini. Legacy models such as GPT-3.5 often display abrupt drops in performance, e.g., in summarization using GlobEnc, necessitating greater caution when reducing tokens.
6.5.  Random & Bottom $k$ tokens
@@ 7.  Conclusion
As our token reduction pipeline heavily relies on attribution methods, misalignment between salience and true causal importance could produce suboptimal performance. The task specificity of the attribution methods also makes them unsuitable for certain tasks, such as reasoning, and evaluation in multi-task settings.
Ethical Consideration
As identified in our work, performance retention in certain tasks might stem from the model’s exposure to benchmark data during pretraining. This raises concerns about the validity of LLM evaluation and the need for contamination-free datasets. Prompt compression can also potentially remove safety-critical tokens and hence should be used with caution.
Bibliographical References
@@
  • Abnar and Zuidema (2020) Samira Abnar and Willem Zuidema. 2020. Quantifying attention flow in transformers. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 4190–4197, Online. Association for Computational Linguistics.
  • Ahmed et al. (2017) Hadeer Ahmed, Issa Traore, and Sherif Saad. 2017. Detection of online fake news using n-gram analysis and machine learning techniques. In International conference on intelligent, secure, and dependable systems in distributed and cloud environments, pages 127–138. Springer.
  • Ainslie et al. (2023) Joshua Ainslie, James Lee-Thorp, Michiel de Jong, Yury Zemlyanskiy, Federico Lebron, and Sumit Sanghai. 2023. GQA: Training generalized multi-query transformer models from multi-head checkpoints. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 4895–4901, Singapore. Association for Computational Linguistics.
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems, 33:1877–1901.
  • Chang et al. (2024) Kaiyan Chang, Songcheng Xu, Chenglong Wang, Yingfeng Luo, Tong Xiao, and Jingbo Zhu. 2024. Efficient prompting methods for large language models: A survey. ArXiv, abs/2404.01077.
  • Chiang et al. (2023) Wei-Lin Chiang, Zhuohan Li, Zi Lin, Ying Sheng, Zhanghao Wu, Hao Zhang, Lianmin Zheng, Siyuan Zhuang, Yonghao Zhuang, Joseph E. Gonzalez, Ion Stoica, and Eric P. Xing. 2023. Vicuna: An open-source chatbot impressing gpt-4 with 90%* chatgpt quality. Accessed: 20-May-25.
  • Child et al. (2019) Rewon Child, Scott Gray, Alec Radford, and Ilya Sutskever. 2019. Generating long sequences with sparse transformers. ArXiv, abs/1904.10509.
  • Chuang et al. (2024) Yu-Neng Chuang, Tianwei Xing, Chia yuan Chang, Zirui Liu, Xun Chen, and Xia Hu. 2024. Learning to compress prompt in natural language formats. In North American Chapter of the Association for Computational Linguistics.
  • Clark (2019) Kevin Clark. 2019. What does bert look at? an analysis of bert’s attention. arXiv preprint arXiv:1906.04341.
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, et al. 2021. Training verifiers to solve math word problems. arXiv preprint arXiv:2110.14168.
  • Dettmers et al. (2023a) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2023a. QLoRA: Efficient finetuning of quantized LLMs. In Thirty-seventh Conference on Neural Information Processing Systems.
  • Dettmers et al. (2023b) Tim Dettmers, Ruslan Svirschevski, Vage Egiazarian, Denis Kuznedelev, Elias Frantar, Saleh Ashkboos, Alexander Borzunov, Torsten Hoefler, and Dan Alistarh. 2023b. Spqr: A sparse-quantized representation for near-lossless llm weight compression.
  • Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. Bert: Pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 conference of the North American chapter of the association for computational linguistics: human language technologies, volume 1 (long and short papers), pages 4171–4186.
  • Fedus et al. (2022) William Fedus, Barret Zoph, and Noam Shazeer. 2022. Switch transformers: Scaling to trillion parameter models with simple and efficient sparsity. Journal of Machine Learning Research, 23(120):1–39.
  • Google DeepMind (2025) Google DeepMind. 2025. Gemini 2.5: Pushing the frontier with advanced reasoning, multimodality, long context, and next generation agentic capabilities. Technical Report. Citation for the Gemini 2.0 Flash Thinking family of models.
  • Gu and Dao (2023) Albert Gu and Tri Dao. 2023. Mamba: Linear-time sequence modeling with selective state spaces. ArXiv, abs/2312.00752.
  • Htut et al. (2019) Phu Mon Htut, Jason Phang, Shikha Bordia, and Samuel R Bowman. 2019. Do attention heads in bert track syntactic dependencies? arXiv preprint arXiv:1911.12246.
  • Hu et al. (2022) Edward J Hu, yelong shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2022. LoRA: Low-rank adaptation of large language models. In International Conference on Learning Representations.
  • Huang et al. (2019) Lifu Huang, Ronan Le Bras, Chandra Bhagavatula, and Yejin Choi. 2019. Cosmos QA: Machine reading comprehension with contextual commonsense reasoning. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 2391–2401, Hong Kong, China. Association for Computational Linguistics.
  • Hudson et al. (2023) Nathaniel C Hudson, J Gregory Pauloski, Matt Baughman, Alok Kamatar, Mansi Sakarvadia, Logan Ward, Ryan Chard, André Bauer, Maksim Levental, Wenyi Wang, et al. 2023. Trillion parameter ai serving infrastructure for scientific discovery: A survey and vision. In Proceedings of the IEEE/ACM 10th International Conference on Big Data Computing, Applications and Technologies, pages 1–10.
  • Islam et al. (2024) Md Ashraful Islam, Mohammed Eunus Ali, and Md Rizwan Parvez. 2024. Mapcoder: Multi-agent code generation for competitive problem solving. arXiv preprint arXiv:2405.11403.
  • Jain and Wallace (2019) Sarthak Jain and Byron C Wallace. 2019. Attention is not explanation. arXiv preprint arXiv:1902.10186.
  • Jang et al. (2024) Lawrence Jang, Yinheng Li, Dan Zhao, Charles Ding, Justin Lin, Paul Pu Liang, Rogerio Bonatti, and Kazuhito Koishida. 2024. Videowebarena: Evaluating long context multimodal agents with video understanding web tasks. arXiv preprint arXiv:2410.19100.
  • Jiang et al. (2023) Huiqiang Jiang, Qianhui Wu, Chin-Yew Lin, Yuqing Yang, and Lili Qiu. 2023. Llmlingua: Compressing prompts for accelerated inference of large language models. In Conference on Empirical Methods in Natural Language Processing.
  • Kaplan et al. (2020) Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. 2020. Scaling laws for neural language models. arXiv preprint arXiv:2001.08361.
  • Kobayashi et al. (2020) Goro Kobayashi, Tatsuki Kuribayashi, Sho Yokoi, and Kentaro Inui. 2020. Attention is not only a weight: Analyzing transformers with vector norms. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 7057–7075, Online. Association for Computational Linguistics.
  • Li and Flanigan (2024) Changmao Li and Jeffrey Flanigan. 2024. Task contamination: Language models may not be few-shot anymore. In Proceedings of the AAAI Conference on Artificial Intelligence, volume 38, pages 18471–18480.
  • Li et al. (2023a) Yixiao Li, Yifan Yu, Qingru Zhang, Chen Liang, Pengcheng He, Weizhu Chen, and Tuo Zhao. 2023a. Losparse: structured compression of large language models based on low-rank and sparse approximation. In Proceedings of the 40th International Conference on Machine Learning, ICML’23. JMLR.org.
  • Li et al. (2023b) Yucheng Li, Bo Dong, Chenghua Lin, and Frank Guerin. 2023b. Compressing context to enhance inference efficiency of large language models. In Conference on Empirical Methods in Natural Language Processing.
  • Liu et al. (2023) Hong Liu, Zhiyuan Li, David Hall, Percy Liang, and Tengyu Ma. 2023. Sophia: A scalable stochastic second-order optimizer for language model pre-training. arXiv preprint arXiv:2305.14342.
  • Ma et al. (2023) Xinyin Ma, Gongfan Fang, and Xinchao Wang. 2023. Llm-pruner: On the structural pruning of large language models. In Advances in Neural Information Processing Systems.
  • Maas et al. (2011) Andrew L. Maas, Raymond E. Daly, Peter T. Pham, Dan Huang, Andrew Y. Ng, and Christopher Potts. 2011. Learning word vectors for sentiment analysis. In Proceedings of the 49th Annual Meeting of the Association for Computational Linguistics: Human Language Technologies, pages 142–150, Portland, Oregon, USA. Association for Computational Linguistics.
  • Meta (2024) Meta. 2024. Introducing meta llama 3: The next generation of our open-source large language model. Meta AI Blog Post and Technical Resource. Citation for the LLaMA-3 8B and LLaMA-3 70B models.
  • Modarressi et al. (2023) Ali Modarressi, Mohsen Fayyaz, Ehsan Aghazadeh, Yadollah Yaghoobzadeh, and Mohammad Taher Pilehvar. 2023. DecompX: Explaining transformers decisions by propagating token decomposition. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 2649–2664, Toronto, Canada. Association for Computational Linguistics.
  • Modarressi et al. (2022) Ali Modarressi, Mohsen Fayyaz, Yadollah Yaghoobzadeh, and Mohammad Taher Pilehvar. 2022. GlobEnc: Quantifying global token attribution by incorporating the whole encoder layer in transformers. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 258–271, Seattle, United States. Association for Computational Linguistics.
  • OpenAI (2025) OpenAI. 2025. OpenAI o3-mini system card. System Card. Citation for the o3-mini model.
  • Radford et al. (2019) Alec Radford, Jeff Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever. 2019. Language models are unsupervised multitask learners.
  • Santra et al. (2023) Bishal Santra, Sakya Basak, Abhinandan De, Manish Gupta, and Pawan Goyal. 2023. Frugal prompting for dialog models. arXiv preprint arXiv:2305.14919.
  • Schmidgall et al. (2025) Samuel Schmidgall, Yusheng Su, Ze Wang, Ximeng Sun, Jialian Wu, Xiaodong Yu, Jiang Liu, Michael Moor, Zicheng Liu, and Emad Barsoum. 2025. Agent laboratory: Using llm agents as research assistants. arXiv preprint arXiv:2501.04227.
  • Sennrich et al. (2016) Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016. Neural machine translation of rare words with subword units.
  • Shubert et al. (1995) Serena K Shubert, Jan H Spyridakis, Heather K Holmback, and Mary B Coney. 1995. The comprehensibility of simplified english in procedures. Journal of technical writing and communication, 25(4):347–369.
  • Stojkovic et al. (2024) Jovan Stojkovic, Esha Choukse, Chaojie Zhang, Inigo Goiri, and Josep Torrellas. 2024. Towards greener llms: Bringing energy-efficiency to the forefront of llm inference. arXiv preprint arXiv:2403.20306.
  • Tay et al. (2022) Yi Tay, Mostafa Dehghani, Dara Bahri, and Donald Metzler. 2022. Efficient transformers: A survey. ACM Comput. Surv., 55(6).
  • Timiryasov and Tastet (2023) Inar Timiryasov and Jean-Loup Tastet. 2023. Baby llama: knowledge distillation from an ensemble of teachers trained on a small dataset with no performance penalty. In Proceedings of the BabyLM Challenge at the 27th Conference on Computational Natural Language Learning, pages 279–289, Singapore. Association for Computational Linguistics.
  • Wan et al. (2023) Zhongwei Wan, Xin Wang, Che Liu, Samiul Alam, Yu Zheng, et al. 2023. Efficient large language models: A survey. arXiv preprint arXiv:2312.03863, 1.
  • Wit and Gillette (1999) EC Wit and Marie Gillette. 1999. What is linguistic redundancy. University of Chicago.
  • Wu et al. (2025) Yangzhen Wu, Zhiqing Sun, Shanda Li, Sean Welleck, and Yiming Yang. 2025. Inference scaling laws: An empirical analysis of compute-optimal inference for llm problem-solving. In The Thirteenth International Conference on Learning Representations.
  • Yang et al. (2024) John Yang, Carlos E Jimenez, Alexander Wettig, Kilian Lieret, Shunyu Yao, Karthik Narasimhan, and Ofir Press. 2024. Swe-agent: Agent-computer interfaces enable automated software engineering. Advances in Neural Information Processing Systems, 37:50528–50652.
  • Yao et al. (2024) Yiqun Yao, Zheng Zhang, Jing Li, and Yequan Wang. 2024. Masked structural growth for 2x faster language model pre-training. In The Twelfth International Conference on Learning Representations.
  • Zhang et al. (2025) Zheng Zhang, Jinyi Li, Yihuai Lan, Xiangwei Wang, and Hao Wang. 2025. An empirical study on prompt compression for large language models.
  • Zhao et al. (2023a) Wenbo Zhao, Arpit Gupta, Tagyoung Chung, and Jing Huang. 2023a. SPC: Soft prompt construction for cross domain generalization. In Proceedings of the 8th Workshop on Representation Learning for NLP (RepL4NLP 2023), pages 118–130, Toronto, Canada. Association for Computational Linguistics.
  • Zhao et al. (2023b) Wenbo Zhao, Arpit Gupta, Tagyoung Chung, and Jing Huang. 2023b. Spc: Soft prompt construction for cross domain generalization. In Workshop on Representation Learning for NLP.
  • Zhou et al. (2024) Zixuan Zhou, Xuefei Ning, Ke Hong, Tianyu Fu, Jiaming Xu, Shiyao Li, Yuming Lou, Luning Wang, Zhihang Yuan, Xiuhong Li, et al. 2024. A survey on efficient inference for large language models. arXiv preprint arXiv:2404.14294.

### v3

@@ 4.5 Evaluation Metrics
Attribution Type Token Attribution Method $k\%$ Tasks
@@
5 Result and Analysis
5.1 Performance across Tasks
Table 1 reveals that performance degradation becomes more pronounced as the retention threshold decreases, particularly below $60\%$ token retention, with distinct patterns emerging across tasks. For instance, at $50\%$ retention, sentiment classification accuracy drops by $5$-$10\%$ and commonsense QA accuracy by $10$-$15\%$, while summarization metrics like ROUGE-1 decline by $20$-$25\%$, as seen in Figure 2,3. The performance decline is noticeably smaller for higher values of $k$, e.g.,, GlobEnc using $80\%$ tokens achieves close to the baseline performance in classification and QA tasks.
Mathematical reasoning exhibits the steepest fall, with pass@1 scores plummeting to below $15\%$ at lower thresholds, reflecting the loss of critical numerical and logical connectors. We hypothesize that this degradation is more pronounced for reasoning tasks due to their reliance on the sequential integrity of tokens and numerical data. Low $k\%$ thresholds exacerbate this by discarding essential context, amplifying errors in tasks requiring unbroken symbolic chains, while simpler classification and summarization tasks generally tolerate greater omission by inferring from salient cues.
@@ 5.1 Performance across Tasks
5.2 Choice of Attribution Method
Following both Table 1 and Figure 2, for classification and QA tasks, GlobEnc outperforms DecompX across all metrics. However, in reasoning, DecompX outshines GlobEnc. DecompX’s subword-level attribution that includes linear activations may help retain more important tokens for reasoning, but might overwhelm the models in other tasks. For summarization (Figure 3), GlobEnc usually outperforms DecompX, except for the GPT-3.5 model. While DecompX is reported to produce better attribution scores Modarressi et al. (2023), most models find the token sequence filtered by word-level attention, like GlobEnc, to be more useful in retaining performance across most tasks.
5.3 Performance vs. Cost
The cost per million tokens for each model was taken from the respective provider’s website, at the time of writing, and is reported in Table 3. Note that the price is subject to change. The cost of the open-source Llama models was estimated using the lowest available rate on OpenRouter.2
Figures 2 and 3 show the performance difference between the reduced set of tokens and $100\%$ tokens for all the models. For sentiment analysis and question answering, we observe that using a reduced set of tokens generally maintains performance and shows a gentler decline as we decrease the cost. However, reducing the tokens sharply declines the performance in reasoning tasks, making it unsuitable for token reduction. The results are mixed for summarization tasks: while most models have a softer performance decline across the metrics, methods such as GPT-3.5 using GlobEnc experience rapid performance degradation.
5.4 Performance across Models
Figures 2 and 3 illustrate that for smaller models, token attribution offers negligible cost savings relative to the sharper performance degradation it entails. Consequently, token reduction is more practical for larger, costlier models such as o3-mini. Legacy models such as GPT-3.5 often display abrupt drops in performance, e.g., in summarization using GlobEnc, necessitating greater caution when reducing tokens.
5.5 Random & Bottom $k$ tokens
@@ 6 Conclusion
As our token reduction pipeline heavily relies on attribution methods, misalignment between salience and true causal importance could produce suboptimal performance. The task specificity of the attribution methods also makes them unsuitable for certain tasks, such as reasoning and evaluation in multi-task settings.
We also present a few ethical considerations. As identified in our work, performance retention in certain tasks might stem from the model’s exposure to benchmark data during pretraining. This raises concerns about the validity of LLM evaluation and the need for contamination-free datasets. Prompt compression can also potentially remove safety-critical tokens and hence should be used with caution.
Appendix A Preliminaries: Token Attribution
A.1 GlobEnc
@@
$\hat{\mathbf{A}}_{l}$ $=0.5\bar{\mathbf{A}}_{l}+0.5\mathbf{I}$ (12)
A.2 DecompX
where the normalized contribution is given by:
$\mathbf{z}_{\text{FFN}}^{\ell}$ $=\text{FFN}(\tilde{\mathbf{z}}_{i}^{\ell})$ (15)
$y_{c}=\sum_{k=1}^{N}y_{c\Leftarrow k}$ (18)
Appendix B Proof
We provide the derivation of Theorem 1.
Proof.
We order $D=\{d_{1},\dots,d_{q}\}$ and let $A_{r}:=\{d_{1},\dots,d_{r}\}$ be the set of tokens deleted after $r$ steps and $A_{0}=\emptyset$. For each $r$ and $u\in A_{r-1}$,
$\Delta_{d_{r},u}f(A_{r-1}\setminus\{u\})\leq\gamma_{\tau}$ [Eq. 5]
$\implies$ $\Delta_{d_{r}}\Delta_{u}f(A_{r-1}\setminus\{u\})\;\leq\;\gamma_{\tau}$ [Eq. 3]
$\implies$ $\Delta_{d_{r}}f(A_{r-1})-\Delta_{d_{r}}f(A_{r-1}\setminus\{u\})\leq+\gamma_{\tau}$
$\implies$ $\Delta_{d_{r}}f(A_{r-1})\leq\Delta_{d_{r}}f(A_{r-1}\setminus\{u\})+\gamma_{\tau}.$
Iterating over all $u\in A_{r-1}$ gives
$\Delta_{d_{r}}f(A_{r-1})$ $\leq\Delta_{d_{r}}f(\emptyset)+\gamma_{\tau}(r-1)$
$\leq f(\{d_{r}\})-f(\emptyset)+\gamma_{\tau}(r-1)$
$\leq f(\{d_{r}\})+\gamma_{\tau}(r-1)$
[…]
Formulating the telescopic sum:
$f(D)=f(A_{q})$ $=\sum_{r=1}^{q}(f(A_{r})-f(A_{r-1}))$
$=\sum_{r=1}^{q}\Delta_{d_{r}}f(A_{r-1})$ [Eq. 2]
$\leq\sum_{r=1}^{q}C_{\tau}s_{d_{r}}+\gamma_{\tau}\sum_{r=1}^{q}(r-1).$
$\text{Hence, }f(D)$ $\leq C_{\tau}\sum_{i\in D}s_{i}+\frac{\gamma_{\tau}}{2}q(q-1).$$\qed$